In a survey, 28 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $48 and standard deviation of $3. Find the margin of error at a 80% confidence level.

Answer by rothauserc(4718)

standard error(SE) = 3 / square root(28) = 0.5669
:
alpha(a) = 1 - (80/100) = 0.20
p* = 1 - a/2 = 0.90
:
critical value(CV) is determined by finding in the z-tables the z-score associated with p*, CV = 1.28
:
margin of error(ME) = CV * SE = 1.28 * 0.5669 = 0.7256
: